# Quantum Wave Function Collapse for Procedural Content Generation

### v1

@@ 1.1 PIPCG
represents the generated partial content instance at the end of iteration $k$ with $|C(k)|=k$, $C(k)=C(k-1)\cup\{(s(k),v_{s(k)})\}$, $s(i)\neq s(j)\,\forall\,i,j\in[1,k]\land i\neq j$. Here, $\mathcal{C}(k)$ denotes all possible partial content instances at the end of iteration $k$. Thus, $|\mathcal{C}(k)|=\sum_{I\in J(k)}\prod_{i\in I}|A_{i}|$, where $J(k):=\{I\,|\,I\in[1,N]\land|I|=k\}$ with $|J(k)|={N\choose k}$. It is convenient to define $C(0):=\{\}$ and $\mathcal{C}(0):=\{C(0)\}$. In addition,
$S^{+}(C(l)):=\{s(1),\dots,s(l)\}$ (3)
denotes the set of segment indices (dropping the values) that are contained in the partial content instance $C(l)$ with $|S^{+}(C(l))|=l$ and
$S^{-}(C(l)):=[1,N]\setminus S^{+}(C(l))$ (4)
denotes the remaining segment indices that are not contained in $C(l)$ with $|S^{-}(C(l))|=N-l$ for all $C(l)\in\mathcal{C}(l)$ and for all $l\in[0,N]$.
After $N$ iterations, the newly generated content instance is complete, i.e., $C(N)\equiv C$ and $\mathcal{C}(N)\equiv\mathcal{C}$. In the following, the three steps within each iteration $k$ are explained in more detail.
1.1.2 First step.
@@
For the first step, an identifier is selected by drawing a sample $s(k)$ from the random variable
@@
The probability distribution $p_{k}^{\mathrm{s}}(i|C^{\prime})$ with support $i\in[1,N]$ and $C^{\prime}\in\mathcal{C}(k-1)$ is a user-defined parameter of the procedure that determines how new segment indices $l$ are selected based on the partial content instance $C^{\prime}=C(k-1)$ that has already been generated up to the previous iteration $k-1$. The dependence on all previously chosen segments makes the process generally non-Markovian. It is required that two conditions hold true for the probability distribution $p_{k}^{\mathrm{s}}(i|C^{\prime})$:
  1. S1
@@
The segment identifier selection mechanism is one of the main ingredients of CWFC. Here, it is presented in very generic way. The common entropic approach will be explained further below along with possible variants that are more suitable for a quantum approach.
1.1.3 Second step.
After the identifier $s(k)$ has been chosen, the second step is to select a corresponding value $v_{s(k)}$ for the segment. For this purpose, a sample $v_{s(k)}$ is drawn from the random variable
@@
The probability distribution $p^{\mathrm{v}}_{k}(\nu|i,C^{\prime})$ with support $\nu\in\cup_{A\in\mathcal{A}}A$, $i\in[1,N]$ and $C^{\prime}\in\mathcal{C}(k-1)$ is a user-defined parameter of the procedure that determines how new values are selected based on the selected identifier $i=s(k)$ of the current iteration and the partial content instance $C^{\prime}=C(k-1)$ that has already been generated up to the previous iteration $k-1$. It is required that two conditions hold true for the probability distribution $p^{\mathrm{v}}_{k}(\nu|i,C^{\prime})$:
  1. V1
    Only values from the corresponding alphabet are available, i.e., $p^{\mathrm{v}}_{l}(\nu=v|i=s,C^{\prime}=C(l))=0\,\forall\,v\not\in A_{s}\,\forall\,s\in[1,N]\,\forall\,C(l)\in\mathcal{C}(l-1)\,\forall\,l\in[1,N]$.
  2. V2
    At least one value from the corresponding alphabet has to be available, i.e., $\forall\,l\in[1,N]\,\forall\,C(l-1)\in\mathcal{C}(l-1)\forall s\in S^{-}(C(l-1))\exists v\in A_{s}:p_{l}^{\mathrm{v}}(\nu=v|i=s,C^{\prime}=C(l))>0$.
As explained further below, it is also possible to specify $p^{\mathrm{v}}_{k}(\nu|i,C^{\prime})$ implicitly based on a set of rules, which is the default approach for CWFC and much more convenient for practical purposes than an explicit definition. To simplify the notation a joint distribution for the segment identifier and value selection $p_{k}(i,\nu|C^{\prime}):=p_{k}^{\mathrm{s}}(i|C^{\prime})p_{k}^{\mathrm{v}}(\nu|i,C^{\prime})$ can be defined.
1.1.4 Third step.
In the third step, the selected tuple $(s(k),v_{s(k)})$ that constitutes the newly generated segment is added to the already generated segments of the content instance, i.e., $C(k)=C(k-1)\cup\{(s(k),v_{s(k)})\}$. If $k<N$, the process is repeated and another segment is generated for the next iteration $k+1$. If $k=N$, the PIPCG is finished.
1.1.5 Instance distribution.
An essential concept of PCG is, that only a limited subset of all possible content instances $C\in\mathcal{C}$ is generated according to the designer’s aesthetic preferences, a given rule set, or some other underlying logic. Otherwise, the instances could be drawn directly from $\mathcal{C}$ with much less effort. For the presented formalism, the probability to generate a partial content instance $C(k)$ in the form of eq. 2 is given by
$p(C(k)):=\sum_{\sigma\in S_{k}(C(k))}\prod_{l=1}^{k}q_{l}^{k}(\sigma,C(k))$ (7)
with the abbreviation
@@
Here, $S_{k}(C(k))$ denotes the set of all permutations of $S^{+}(C(k))$, eq. 3, with $|S_{k}(C(k))|=k!$ and $\sigma_{l}$ represents the $l$th value of a permutation $\sigma\in S_{k}(C(k))$ for $l\in[1,k]$. The notation in section 1.1.5 is to be understood in such a way that $C^{\prime}=\{\}$ for $l=1$. Moreover, $v_{\sigma_{m}}$ in section 1.1.5 represents the value of the segment with identifier $\sigma_{m}$ from the content instance $C(k)$ for $m\in[1,l]$. By definition, the partial content instance $C(k)$ can only be generated if $p(C(k))>0$. In other words, the support
$\underline{\mathcal{C}}(k):=\{C(k)\,|\,C(k)\in\mathcal{C}(k)\,\land\,p(C(k))>0\}$ (9)
of $p(C(k))$ contains all partial content instances that can be generated for the given $k$.
For $k=N$, consider a complete content instance $C$ in the form of eq. 1 with values $v_{1},\dots,v_{N}$. In this case, eq. 7 reads
$p(C)=p(v_{1},\dots,v_{N}):=\sum_{\sigma\in S_{N}}\prod_{l=1}^{N}q_{l}^{N}(\sigma,C),$ (10)
where $S_{N}:=S_{N}(C)$ is the set of all permutations of $[1,N]$. In fact, performing the presented PIPCG procedure corresponds to drawing a sample from this distribution such that the content instance can be considered as a random variable $\boldsymbol{C}\sim p(C)$. Consequently, $\underline{\mathcal{C}}(N)=\underline{\mathcal{C}}:=\{C\,|\,C\in\mathcal{C}\,\land\,p(C)>0\}$ denotes the set of all content instances that can be generated.
1.2 Classical Wave Function Collapse (CWFC)
CWFC is a special case of the previously described PIPCG. In this manuscript, only a simplified form of the simple tiled model of CWFC is considered, as it is sufficient to capture the key components of the approach. A detailed description of the original CWFC implementation can be found in, e.g., Refs. [2, 3].
For the common CWFC approach, a single alphabet is used for all segments, i.e., $A_{i}=A\,\forall\,i\in[1,N]$. Given $W$ symbols, one can presume without loss of generality that
$A:=[1,W].$ (11)
Furthermore, the segment identifier selection is performed with an entropic approach, whereas the value selection is pattern-based. These two concepts will be explained in more detail in the following.
1.2.1 Value selection.
For the pattern-based value selection, the segments are organized in such a way that they have a predefined adjacency relationship to each other that does not change. For example, consider that the pixels of an image constitute the segments. Choosing adjacency between the nearest neighbors leads to an adjacency configuration with four directions (right, up, left, down).
Formally, each direction within a adjacency configuration is specified by an index $d\in[1,D]$, where $D$ denotes the total number of directions. The direction-based adjacency relationship between different segments is then defined by the coefficients $\alpha^{d}_{i,j}\in\mathbb{B}:=\{0,1\}$ for $i,j\in[1,N]$ and $d\in[1,D]$. If $\alpha^{d}_{i,j}=1$, segment $i$ is connected to segment $j$ in direction $d$; otherwise, it is not connected. For each direction $d$, these coefficients form an adjacency matrix $\alpha^{d}\in\mathcal{B}:=\mathbb{B}^{N\times N}$ of a directed graph with $N$ vertices that represent the segments. Hence, the segment adjacency is in fact an abstract concept that is not necessarily related to the visual form of the content. An example is shown in fig. 1.
Based on the chosen adjacency configuration, a pattern-based value selection can be defined. For this purpose, a more general rule-based selection is introduced first and then the pattern-based selection is presented as a special case.
For a rule-based value selection, consider a set of $m$ rules
$R:=\{r_{1},\dots,r_{m}\}\in\mathcal{R},$ (12)
from the set of of all possible rulesets $\mathcal{R}$, where each rule
$r_{i}:=r_{i}(v,w):=(v,w)$ (13)
for $i\in[1,m]$ consists of a value $v\in A$ and a weight function $w:=w(i,\alpha^{d},C^{\prime})\in\mathbb{R}_{\geq 0}$ with
$w:[1,N]\times\mathcal{B}\times\overline{\mathcal{C}}\rightarrow\mathbb{R}_{\geq 0},$ (14)
where
$\overline{\mathcal{C}}:=\cup_{k=0}^{N-1}\mathcal{C}(k)$ (15)
denotes the unified set of all partial contents. Given a ruleset $R\in\mathcal{R}$, the rule-based version of the probability distribution from eq. 6 is defined as
$p_{k}^{\mathrm{v}}(\nu|i,C^{\prime})=p_{k}^{\mathrm{v,R}}(\nu|i,C^{\prime}):=\frac{F_{k}(\nu,i,C^{\prime})}{\overline{F}_{k}(i,C^{\prime})}$ (16)
with the abbreviations
$F_{k}(\nu,i,C^{\prime}):=\sum_{\begin{subarray}{c}(v,w)\in R\\ v=\nu\end{subarray}}w(i,\alpha^{d},C^{\prime})$ (17)
and
$\overline{F}_{k}(i,C^{\prime}):=\sum_{\nu\in A}F_{k}(\nu,i,C^{\prime}),$ (18)
respectively, for $\nu\in A$, $i\in[1,N]$, $C^{\prime}\in\mathcal{C}(k)$ for $k\in[1,N]$, and $\alpha^{d}\in\mathcal{B}$ for $d\in[1,D]$. Summarized, the set of rules $R$ determines how values are selected based on the non-vanishing results of the respective weight functions. Note that condition V1 is always satisfied, whereas in case of $\overline{F}_{k}(i,C^{\prime})\overset{!}{=}0$ for any $C^{\prime}\in\mathcal{C}(k)$ with $p_{k}^{\mathrm{s,H}}(i|C^{\prime})p(C^{\prime})>0$, condition V2 is violated.
Pattern-based value selection is a special case of rule-based value selection, where the weight function
$w(i,\alpha^{d},C^{\prime})\!:=\!w^{\mathrm{P}}(i,\alpha^{d},C^{\prime},P,u)\!:=\!u\tau(i,\alpha^{d},C^{\prime},P)$ (19)
of each rule
$r_{i}:=r_{i}^{\mathrm{P}}(v,P,u):=(v,w^{\mathrm{P}}(i,\alpha^{d},C^{\prime},P,u))$ (20)
for $i\in[1,m]$ is defined by a non-vanishing factor $u\in\mathbb{R}_{>0}$ and a pattern
$P:=\{(d_{1},v_{1}),\dots,(d_{n},v_{n})\}\in\mathcal{P}$ (21)
consisting of a set of $n\leq D$ direction-value pairs with $d_{i}\in[1,D]$ for $i\in[1,n]$, $d_{i}\neq d_{j}\,\forall\,i,j\in[1,n]\land i\neq j$, and $v_{i}\in A$ for $i\in[1,n]$, where
$\mathcal{P}:=\{P\,|\,P\in([1,D]\times A)^{n}\,\forall\,n\in[0,D]\}$ (22)
denotes the set of all possible patterns. A pattern $P$ is fulfilled, if
$\tau(i,\alpha^{d},C^{\prime},P):=\prod_{(d,v)\in P}\prod_{\begin{subarray}{c}(s,v^{\prime})\in C^{\prime}\\ \alpha^{d}_{i,s}=1\end{subarray}}\delta_{vv^{\prime}}\in[0,1]$ (23)
evaluates to one. Here, $\delta$ represents the Kronecker delta. The set of all possible pattern-based rulesets is denoted by $\mathcal{R}^{\mathrm{P}}\subset\mathcal{R}$.
Summarized, the weight function, eq. 19, yields $u$ if the pattern $P$ applies and zero otherwise. In line with expectations, only adjacent segments (i.e., $\alpha^{d}_{i,s}=1$) have an influence on eq. 16 according to eq. 23.
Furthermore, the rules from eq. 20 can be rewritten as
$r_{i}^{\mathrm{P}}(v,P,u)=(v,u,P)$ (24)
for $i\in[1,m]$ with a value $v\in A$, a factor $u\in\mathbb{R}_{>0}$, and a pattern $P\in\mathcal{P}$. For the pattern-based value selection, eq. 16 can be reformulated as
$p_{k}^{\mathrm{v,R}}(\nu|i,C^{\prime})$ $=p_{k}^{\mathrm{v,P}}(\nu|i,C^{\prime})$
$:=\sum_{\begin{subarray}{c}(v,u,P)\in R\\ v=\nu\end{subarray}}\gamma(u,i,C^{\prime})\tau(i,\alpha^{d},C^{\prime},P)$ (27)
with the coefficients
$\gamma(u,i,C^{\prime}):=\frac{u}{\overline{F}_{k}(i,C^{\prime})}\in[0,1]$ (28)
that correspond to the normalized weights of the rules.
In the context of this manuscript, one variant (that lies between rule- and pattern-based selection) is particularly important, namely the use of a functional factor instead of a constant factor for pattern-based rules of the form of eq. 24. Specifically, a factor
$u:=u(i,C):[1,N]\times\overline{\mathcal{C}}\rightarrow\mathbb{R}$ (29)
is considered that can depend both on the coordinate $i\in[1,N]$ and/or the already generated content $C\in\overline{\mathcal{C}}$, eq. 15. By mapping to zero, this functional factor can disable a pattern for certain coordinates or partial contents.
1.2.2 Identifier selection.
For the entropic segment identifier selection, the probability distribution from eq. 5 is defined as
@@ 1.2 Classical Wave Function Collapse (CWFC)
with the set of optimal solutions
$I_{k}(C^{\prime}):=\{i\,|\,i\in S^{-}(C^{\prime})\land H_{k}(i,C^{\prime})=h_{k}(C^{\prime})\}$ (31)
of the optimization task
$h_{k}(C^{\prime}):=\operatorname*{arg\,min}_{j\in S^{-}(C^{\prime})}H_{k}(j,C^{\prime}),$ (32)
where
$H_{k}(i,C^{\prime}):=-\sum_{v\in A}$ $p_{k}^{\mathrm{v,P}}(\nu=v|i,C^{\prime})$
×$\times\ln p_{k}^{\mathrm{v,P}}(\nu=v|i,C^{\prime})$ (33)
denotes the Shannon information entropy for $C^{\prime}\in\mathcal{C}(k-1)$, $k\in[1,N]$ That is, the segments with the smallest Shannon entropy with respect to the available value options are chosen.
1.3 CWFC Example: Checkerboard
As a simple example, consider the generation of images with black and white checkerboard patterns. The pixels of an image constitute the segments and the alphabet consists of only $W=2$ symbols standing for the two colors. As an adjacency configuration, the nearest-neighbor adjacency from fig. 1 is presumed.
A checkerboard image can be achieved with only two patterns of the form of eq. 21, namely $P_{1}:=\{(1,1),(2,1),(3,1),(4,1)\}$ and $P_{2}:=\{(1,2),(2,2),(3,2),(4,2)\}$. The corresponding ruleset, eq. 12, reads $R=\{r_{1}^{\mathrm{P}},r_{2}^{\mathrm{P}}\}$ and therefore contains the $m=2$ rules $r_{1}^{\mathrm{P}}(v=2,P=P_{1},u=1)$ and $r_{2}^{\mathrm{P}}(v=1,P=P_{1},u=1)$ based on eq. 20, see fig. 2.
Figure 2: Visualization of pattern-based rules. (a) Visualization scheme. The center tile corresponds to the target segment $i$ of the weight function, eq. 19, with a target value of $v$, whereas the adjacent tiles represent the required values $v_{1},\dots,v_{4}\in\{1,2\}$ of the four adjacent segments in the respective directions $d=1,\dots,d=4$. (b) Pattern-based rules $r_{1}^{\mathrm{P}}(v=2,P=P_{1},u=1)$ and $r_{2}^{\mathrm{P}}(v=1,P=P_{2},u=1)$, eq. 20, that can be used for the generation of images with black and white checkerboard patterns.
The iterative process of CWFC is sketched in fig. 3 for a $3\times 3$ image (represented by $N=9$ segments). Only two content instances can be generated, $C_{1}:=\{(1,1),\dots\}$ and $C_{2}:=\{(1,2),\dots\}$ with equal probability, eq. 10. The choice of the first value $v_{s(1)}$ already determines the choices of all other values.
Figure 3: CWFC for the generation of $3\times 3$ images with checkerboard patterns defined by the rules from fig. 2. In each iteration $k\in[1,9]$, three steps take place: (i) a segment identifier $s(k)$ is drawn, eq. 30, (ii) a corresponding value $v_{s(k)}$ is drawn, section 1.2.1, and (iii) the newly generated identifier-value pair is added as a segment to the content instance $C(k)$. For $s(k)$ and $v(k)$, the uniformly distributed set of possible choices are listed. The choices for $s(k)$ depend on the Shannon entropy $H_{k}$, section 1.2.2, of each undefined segment, as shown on the right. The choices for $v(k)$ depend on the fulfillment of the two rules $r_{1}^{\mathrm{P}}$ and $r_{2}^{\mathrm{P}}$, as listed.
2 QUANTUM METHODS
Based on the previously established classical concepts, a QWFC method and a HWFC method are proposed in the following. A software implementation is available online [9].
2.1 Quantum Wave Function Collapse (QWFC)
So far, the procedural generation concepts have been presented from a purely classical perspective. However, it seems natural that the probability distribution of the content instances $p(C)$, eq. 10, can also be realized with the help of a quantum circuit by utilizing the intrinsic randomness of quantum physics [10]. This is the very idea on which QWFC is based.
A single alphabet $A$ with $W$ symbols is presumed, eq. 11. Furthermore, two procedures control the content generation, namely segment identifier selection and value selection.
2.1.1 Identifier selection.
The Identifier selection is performed with a predefined order given by the vector $\sigma\in S_{N}$, which is the set of all permutations of $[1,N]$. Hence,
$p_{k}^{\mathrm{s}}(i|C^{\prime})=p_{k}^{\mathrm{s,o}}(i|C^{\prime}):=\delta_{i\sigma_{i}}.$ (34)
This approach leads to (partial) content instances $\mathcal{C}(k)=\mathcal{C}_{\sigma}(k)$, eq. 2, of the form
$\mathcal{C}_{\sigma}(k):=\{C\,|\,$ $C=\{(\sigma_{1},v_{\sigma_{1}}),\dots,(\sigma_{k},v_{\sigma_{k}})\}$
$\,\forall\,v_{\sigma_{l}}\in A\,\forall\,l\in[1,k]\}$ (35)
for $k\in[1,N]$ and $\mathcal{C}_{\sigma}(0):=\{\{\}\}$.
2.1.2 Value selection.
The value selection is pattern-based in analogy to section 1.2.1. In particular, the value for each segment is encoded by a set of qubits, which requires
$q:=\lceil\log_{2}W\rceil$ (36)
qubits for each segment $i\in[1,N]$. The group of $q$ qubits that represent the value for segment $i$ is denoted by
$\mathcal{Q}_{i}:=\{q_{i}^{1},\dots,q_{i}^{q}\},$ (37)
where $q_{i}^{j}$ references the $j$th qubit of this group for $j\in[1,q]$. The state of these qubits encodes the value $v_{i}$ of the segment $i$ in binary as the tensor product sequence of qubit states
$\bigotimes_{j=1}^{q}\ket{b_{j}}=\Ket{v_{i}-1=\sum_{j=1}^{q}b_{j}2^{j-1}}_{i}=\ket{v_{i}-1}_{i}\in\mathcal{H}_{i}$ (38)
with $b_{j}\in\mathbb{B}$, $v_{i}\in A$, and $i\in[1,N]$, where $\mathcal{H}_{i}$ denotes the joint Hilbert space of all qubit from $\mathcal{Q}_{i}$.
In total,
$Q:=Nq$ (39)
qubits are required to represent the value of all $N$ segments. The joint Hilbert space of all qubits from $\bigcup_{i=1}^{N}\mathcal{Q}_{i}$ is given by $\mathcal{H}:=\bigotimes_{i=1}^{N}\mathcal{H}_{i}$ and a state |$\ket{\Psi}\in\mathcal{H}$ can realize samples from a specific probability distribution of content instances
$p(C\!=\!\{(1,v_{1}),\dots,(N,v_{N})\})=|\bra{\Psi}\bigotimes_{i=1}^{N}\ket{v_{i}-1}_{i}|^{2}.$ (40)
based on a projective measurement. After these preliminary considerations, the remaining question is how such a state can be prepared constructively with quantum gates from a given pattern-based ruleset. This question will be answered in the following.
2.1.3 Quantum circuit.
Initially, all qubits are prepared in the ground state, i.e., |$\ket{\Psi}=\ket{0}:=\bigotimes_{i=1}^{N}\ket{0}_{i}$. Then, given a pattern-based ruleset $R\in\mathcal{R}^{\mathrm{P}}$, eq. 12, a series of operators is applied, one for each iteration $k\in[1,N]$ such that |$\ket{\Psi}:=U_{N}\cdots U_{1}\ket{0}$ represents the final state after $k$ iterations. In each iteration $k$, the state of the qubits $\mathcal{Q}_{\sigma_{k}}$ (that represent the value for segment $\sigma_{k}$) is prepared conditioned on the state of (a subset of) the qubits $\mathcal{Q}_{\sigma_{1}},\dots,\mathcal{Q}_{\sigma_{k-1}}$ (that represent the values of the already prepared segments $\sigma_{1},\dots,\sigma_{k-1}$), which leads to an entangled joint state. A sketch is shown in fig. 4 (a).
Figure 4: QWFC circuit with $\sigma=(1,\dots,N)$. (a) Circuit layout with an exemplarily alphabet $A$, eq. 11, with $W=16$ symbols, which requires a set $\mathcal{Q}_{i}$, eq. 37, of $q=4$ qubits, eq. 36, for each segment with $i\in[1,N]$. In each iteration $k\in[1,N]$, the operator $U_{k}$, eq. 41, prepares the qubits from $\mathcal{Q}_{k}$ in a superposition state |$\ket{C^{\prime}}_{k}$, eq. 45, entangled with (a subset of) the qubits from $\mathcal{Q}_{1},\dots,\mathcal{Q}_{k-1}$. All qubits are measured to obtain $p(C)$, eq. 40. (b) Gate decomposition of $U_{k}$ into the components of $U_{k}^{\mathrm{coin}}$, eq. 42, which consists of pairs $U_{k}^{\mathrm{ctrl}}(C^{\prime})\otimes U_{k}^{\mathrm{init}}(C^{\prime})$, eqs. 43 and 44. Each of these pairs represent a conditional loading of a probability distribution to realize |$\ket{C^{\prime}}_{k}$ (which is not explicitly shown). The combined control operators act on the qubits from $\mathcal{Q}_{k}^{+}$ eq. 47, whereas the initialization operators act on the qubits from $\mathcal{Q}_{k}$.
Specifically,
@@ 2.1 Quantum Wave Function Collapse (QWFC)
for $k\in[1,N]$ contains $\mathbbm{1}_{i}\in\mathcal{H}_{i}$, the unit operator for $i\in[1,N]$ and two other operators. First, the conditional initialization operator

### v2

@@ CONCLUSION
Presume the generated partial content instance $C(k)\in\mathcal{C}(k)$ for $k\in[1,N]$, where $\mathcal{C}(k)$ denotes all possible partial content instances at the end of iteration $k$ with $\mathcal{C}(N)\equiv\mathcal{C}$. One has $|C(k)|=k$, $C(k)=C(k-1)\cup\{(s(k),v_{s(k)})\}$, $s(i)\neq s(j)\,\forall\,i,j\in[1,k]\land i\neq j$, and $|\mathcal{C}(k)|=\sum_{I\in J(k)}W^{|I|}$, where $J(k):=\{I\,|\,I\in[1,N]\land|I|=k\}$ with $|J(k)|={N\choose k}$. Let $C(0):=\{\}$ and $\mathcal{C}(0):=\{C(0)\}$. Furthermore, $S^{+}(C(l)):=\{s(1),\dots,s(l)\}$ with $|S^{+}(C(l))|=l$ and $S^{-}(C(l)):=[1,N]\setminus S^{+}(C(l))$ with $|S^{-}(C(l))|=N-l$ for all $C(l)\in\mathcal{C}(l)$ and for all $l\in[0,N]$.
@@
In the following, the three steps within each iteration $k$ of PIPCG are explained. In the first step, an identifier is selected by drawing a sample $s(k)$ from the random variable
@@
The probability distribution $p_{k}^{\mathrm{s}}(i|C^{\prime})$ with support $i\in[1,N]$ and $C^{\prime}\in\mathcal{C}(k-1)$ is a user-defined parameter of the procedure that determines how new segment indices $k$ are selected based on the partial content instance $C^{\prime}=C(k-1)$ that has already been generated up to the previous iteration $k-1$ (making PIPCG non-Markovian). Two conditions must hold:
  1. S1)
@@ 2. S2)
In the second step, a value $v_{s(k)}$ is selected for the segment by drawing a sample $v_{s(k)}$ from the random variable
@@
The probability distribution $p^{\mathrm{v}}_{k}(\nu|i,C^{\prime})$ with support $\nu\in A$, $i\in[1,N]$ and $C^{\prime}\in\mathcal{C}(k-1)$ is a user-defined parameter of the procedure that determines how new values are selected based on the selected identifier $i=s(k)$ of the current iteration and $C^{\prime}=C(k-1)$. Two conditions must hold:
  1. V1)
    Only values from the alphabet are available, i.e., $p^{\mathrm{v}}_{l}(\nu=v|i=s,C^{\prime}=C(l))=0\,\forall\,v\not\in A\,\forall\,s\in[1,N]\,\forall\,C(l)\in\mathcal{C}(l-1)\,\forall\,l\in[1,N]$.
  2. V2)
    At least one value from the alphabet has to be available, i.e., $\forall\,l\in[1,N]\,\forall\,C(l-1)\in\mathcal{C}(l-1)\forall s\in S^{-}(C(l-1))\exists v\in A:p_{l}^{\mathrm{v}}(\nu=v|i=s,C^{\prime}=C(l))>0$.
In the third step, $C(k)=C(k-1)\cup\{(s(k),v_{s(k)})\}$. The process is repeated for $k\rightarrow k+1$ until $k=N$.
The probability to generate $C(k)$ is given by $p(C(k)):=\sum_{\sigma\in S_{k}(C(k))}\prod_{l=1}^{k}q_{l}^{k}(\sigma,C(k))$, where
@@ 2. V2)
with $p_{l}(i,\nu|C^{\prime}):=p_{k}^{\mathrm{s}}(i|C^{\prime})p_{k}^{\mathrm{v}}(\nu|i,C^{\prime})$, $S_{k}(C(k))$ as the set of all permutations of $S^{+}(C(k))$ with $|S_{k}(C(k))|=k!$, and the $l$th value of a permutation $\sigma\in S_{k}(C(k))$ denoted by $\sigma_{l}$ for $l\in[1,k]$. In Technical Details of PIPCG, let $C^{\prime}=\{\}$ for $l=1$. $v_{\sigma_{m}}$ represents the value of the segment with identifier $\sigma_{m}$ from the content instance $C(k)$ for $m\in[1,l]$. The support $\underline{\mathcal{C}}(k):=\{C(k)\,|\,C(k)\in\mathcal{C}(k)\land p(C(k))>0\}$ of $p(C(k))$ contains all $C(k)$ that can be generated for the given $k$. For $k=N$, Equation 2 emerges with $p(v_{1},\dots,v_{N}):=\sum_{\sigma\in S_{N}}\prod_{l=1}^{N}q_{l}^{N}(\sigma,C)$ and $\underline{\mathcal{C}}(N)=\underline{\mathcal{C}}$.
Technical Details of CWFC
For the value selection, a general rule-based selection mechanism is introduced first and then the pattern-based selection is presented as a special case. Consider a set of $m$ rules $R\in\mathcal{R}$, Equation 5, with the set of of all possible rulesets $\mathcal{R}$. Each rule $r_{i}:=r_{i}(v,w):=(v,w)$ for $i\in[1,m]$ consists of a value $v\in A$ and a weight function $w:=w(i,\alpha^{d},C^{\prime})\in\mathbb{R}_{\geq 0}$ with $w:[1,N]\times\mathcal{B}\times\overline{\mathcal{C}}\rightarrow\mathbb{R}_{\geq 0}$, where $\overline{\mathcal{C}}:=\cup_{k=0}^{N-1}\mathcal{C}(k)$ and :=$\mathcal{B}:=\{0,1\}^{N\times N}$. Given $R\in\mathcal{R}$, $p_{k}^{\mathrm{v}}(\nu|i,C^{\prime})=p_{k}^{\mathrm{v,R}}(\nu|i,C^{\prime}):=\frac{F_{k}(\nu,i,C^{\prime})}{\overline{F}_{k}(i,C^{\prime})}$, Equation 8, with $F_{k}(\nu,i,C^{\prime}):=\sum_{\begin{subarray}{c}(v,w)\in R\\ v=\nu\end{subarray}}w(i,\alpha^{d},C^{\prime})$ and $\overline{F}_{k}(i,C^{\prime}):=\sum_{\nu\in A}F_{k}(\nu,i,C^{\prime})$ for $\nu\in A$, $i\in[1,N]$, $C^{\prime}\in\mathcal{C}(k)$ for $k\in[1,N]$, and $\alpha^{d}\in\mathcal{B}$ for $d\in[1,D]$. Condition V1 is always satisfied, whereas in case of $\overline{F}_{k}(i,C^{\prime})\overset{!}{=}0$ for any $C^{\prime}\in\mathcal{C}(k)$ with $p_{k}^{\mathrm{s,H}}(i|C^{\prime})p(C^{\prime})>0$, condition V2 is violated.
Pattern-based value selection is a special case of rule-based value selection, where $w(i,\alpha^{d},C^{\prime})\!:=\!w^{\mathrm{P}}(i,\alpha^{d},C^{\prime},P,u)\!:=\!u\tau(i,\alpha^{d},C^{\prime},P)$ of each rule $r_{i}:=r_{i}^{\mathrm{P}}(v,P,u):=(v,w^{\mathrm{P}}(i,\alpha^{d},C^{\prime},P,u))$ for $i\in[1,m]$ is defined by $u\in\mathbb{R}_{>0}$ and $P\in\mathcal{P}$, Equation 3 with $d_{i}\in[1,D]$ for $i\in[1,n]$, $d_{i}\neq d_{j}\,\forall\,i,j\in[1,n]\land i\neq j$, and $v_{i}\in A$ for $i\in[1,n]$. The set of all possible patterns reads $\mathcal{P}:=\{P\,|\,P\in([1,D]\times A)^{n}\,\forall\,n\in[0,D]\}$. A pattern $P$ is fulfilled, if $\tau(i,\alpha^{d},C^{\prime},P):=\prod_{(d,v)\in P}\prod_{(s,v^{\prime})\in C^{\prime}\land\alpha^{d}_{i,s}=1}\delta_{vv^{\prime}}\in[0,1]$ evaluates to one. The set of all possible pattern-based rulesets is denoted by $\mathcal{R}^{\mathrm{P}}\subset\mathcal{R}$. Summarized, the weight function $w(i,\alpha^{d},C^{\prime})$, yields $u$ if the pattern $P$ applies and zero otherwise. Only adjacent segments (i.e., $\alpha^{d}_{i,s}=1$) have an influence on $p_{k}^{\mathrm{v}}(\nu|i,C^{\prime})$.
The rules $r_{i}$ can be rewritten in form of Equation 4. One has $p_{k}^{\mathrm{v,R}}(\nu|i,C^{\prime})=p_{k}^{\mathrm{v,P}}(\nu|i,C^{\prime}):=\sum_{\begin{subarray}{c}(v,u,P)\in R\\ v=\nu\end{subarray}}\gamma(u,i,C^{\prime})\tau(i,\alpha^{d},C^{\prime},P)$ with $\gamma(u,i,C^{\prime}):=\frac{u}{\overline{F}_{k}(i,C^{\prime})}\in[0,1]$. Optionally, a functional factor instead of a constant factor can be used, i.e., $u:=u(i,C):[1,N]\times\overline{\mathcal{C}}\rightarrow\mathbb{R}$.
For the entropic segment identifier selection, one has
@@
in Equation 7 with $I_{k}(C^{\prime}):=\{i\,|\,i\in S^{-}(C^{\prime})\land H_{k}(i,C^{\prime})=h_{k}(C^{\prime})\}$ and $h_{k}(C^{\prime}):=\operatorname*{arg\,min}_{j\in S^{-}(C^{\prime})}H_{k}(j,C^{\prime})$, where $H_{k}(i,C^{\prime}):=-\sum_{v\in A}p_{k}^{\mathrm{v,P}}(\nu=v|i,C^{\prime})\ln p_{k}^{\mathrm{v,P}}(\nu=v|i,C^{\prime})$ denotes the Shannon information entropy for $C^{\prime}\in\mathcal{C}(k-1)$, $k\in[1,N]$.
Technical Details of QWFC
Presume a predefined order of segment indices $\sigma\in S_{N}$. Hence, $p_{k}^{\mathrm{s}}(i|C^{\prime})=p_{k}^{\mathrm{s,o}}(i|C^{\prime}):=\delta_{i\sigma_{i}}$ and $\mathcal{C}(k)=\mathcal{C}_{\sigma}(k):=\{C\,|\,C=\{(\sigma_{1},v_{\sigma_{1}}),\dots,(\sigma_{k},v_{\sigma_{k}})\}\,\forall\,v_{\sigma_{l}}\in A\,\forall\,l\in[1,k]\}$ for $k\in[1,N]$ and $\mathcal{C}_{\sigma}(0):=\{\{\}\}$.
The value selection is pattern-based in analogy to CWFC, which together with the identifier selection leads to Equation 6. Given $\sigma$ and $R\in\mathcal{R}^{\mathrm{P}}$, the task is to prepare |$\ket{\Psi}$ from Equation 6 constructively with quantum gates. Let |$\ket{\Psi}:=U_{N}\cdots U_{1}\ket{0}$ as in Figure 4(a). Then,
@@
for $k\in[1,N]$, which contains $\mathbbm{1}_{i}\in\mathcal{H}_{i}$, the unit operator for $i\in[1,N]$ and two other operators.
First, the conditional initialization operator